solve
16x^4+47x^2=3






Answer by dfrazzetto(283)


16x^4+47x^2=3
16x^4 + 47x^2 - 3 = 0
(16x^2 -1)(x^2 + 3) = 0
16x^2 - 1 = 0; x^2 + 3 = 0
16x^2 = 1; x^2 = -3 (no solution, or radical 3*i, if using complex numbers)
x^2 = 1/16
x = +/- 1/4